# Exploring the Latent Capacity of LLMs for One-Step Text Generation thanks: Under review

### v1

@@ 5 Discussion and Conclusions
We find that both the number and the arrangement of such tokens is crucial for enabling this generation capacity. Interestingly, with only one proto-token, LLMs are unable to generate more than a single token of text. In contrast, two properly arranged proto-tokens can enable the generation of sequences hundreds tokens long. This significant leap in the performance, along the observation that one of the vectors can (in principal) be shared across many texts, suggest that proto-tokens play different functional roles during generation. However, the precise nature of the role differentiation remains unclear.
We find that bigger model size does not universally imply better generation capacity. While larger models in Llama-3 family demonstrate improved reconstruction capacity, Pythia models show no such trend – larger models do not outperform smaller one. Whether this difference is connected to the architectural variations is an open question.
Additionally, we do not observe any consistent relationship between the source of the natural text and the reconstruction ability of LLMs. Surprisingly, even for the texts generated by the LLM itself, the number of successfully reconstructed tokens is the same as for any other natural text. However, for the texts composed of random tokens, performance drops noticeably. This suggests that our reconstruction process does not fully leverage the language modeling capabilities of LLMs, and may instead mostly rely on low-level token patterns.
Although the reconstructed sequences in the non-autoregressive setting are, on average, about two times shorter than those in the autoregressive case, the computational efficiency of single-forward approach allows to achieve up to 279× greater generation throughput.
We also observe that proto-tokens encode more than just the target sequence. Embeddings of the "proto-tokens" corresponding to the different texts generated from the same context are significantly closer to each other than those from unrelated sequences. This indicates that the learned representations capture some potential contextual information.
Finally, we discover that the embedding space in which proto-tokens exist, has very desirable structural properties – proto-tokens corresponding to the same text, form localized and connected regions, enabling smooth transitions via quadratic interpolation. These findings suggest that it may be feasible to build an encoder capable of mapping into this space, opening the door to future work on non-autoregressive inference and representation learning.
@@ 6 Limitations
Although our paper demonstrates the surprising capability of LLMs to generate long sequences in a single forward pass from just two learned embeddings, several important limitations should be acknowledged:
1. Lack of immediate practical application: Most importantly, this work highlights an interesting quirk of LLMs and does not suggest any immediate practical implications or real-life usages for the method.
2. Architectural dependence: The method demonstrates different behavior across model families, suggesting some architectural dependence. As a results, our method may potentially not generalize to other model architectures.

### v2

@@ 5 Discussion and Conclusions
We find that both the number and the arrangement of such tokens is crucial for enabling this generation capacity. Interestingly, with only one proto-token, LLMs are unable to generate more than a single token of text. In contrast, two properly arranged proto-tokens can enable the generation of sequences hundreds of tokens long. This significant leap in the performance, along the observation that one of the vectors can (in principle) be shared across many texts, suggests that proto-tokens play different functional roles during generation.
We find that bigger model size does not universally imply better generation capacity. While larger models in Llama-3 family demonstrate improved reconstruction capacity, Pythia models show no such trend – larger models do not perform better.
Additionally, we do not observe any consistent relationship between the source of the natural text and the reconstruction ability of LLMs. Surprisingly, even for the texts generated by the LLM itself, the number of successfully reconstructed tokens is the same as for any other natural text. However, with random-token sequences, performance drops noticeably. This suggests that our reconstruction process does not fully leverage the language modeling capabilities of LLMs, and may instead mostly rely on low-level token patterns.
Although the reconstructed sequences in the non-autoregressive setting are, on average, two times shorter than those in the autoregressive case, the efficiency of single-forward approach allows to achieve up to 279× greater generation throughput.
We also observe that proto-tokens encode more than just the target sequence. Embeddings of the "proto-tokens" corresponding to the different texts generated from the same context are, on average, closer to each other than those from unrelated sequences. This indicates that learned representations capture some potential contextual information.
Finally, we discover that the embedding space of proto-tokens has very desirable structural properties – proto-tokens corresponding to the same text, form localized and connected regions with smooth transitions via quadratic interpolation. These findings suggest that it may be feasible to build an encoder capable of mapping into this space, opening the door to future work on non-autoregressive inference and representation learning.
We view this work as an existence proof: certain text representations can elicit multi-token behavior in frozen, single-token LLMs. Making them practical requires training an encoder that maps text into these representations. As a future work, we plan to study decoders that either generate a suffix from an encoded prefix or reconstruct the entire text. Depending on the setup, such systems could enable multi-token or chunk-wise generation, learned compression or RAG (potentially by extracting information directly without decoding).
@@ 6 Limitations
1. Lack of immediate practical application: Most importantly, this work highlights an interesting quirk of LLMs and does not suggest any immediate practical implications or real-life usages for the method yet, as direct proto-token optimization should be replaced with parametrized encoder for any practical application.
2. Architectural dependence: The method demonstrates different behavior across model families, suggesting some architectural dependence. As a result, our method may potentially not generalize to other model architectures.
@@
4. Evidence, not bounds: Our 5k-step optimization budget may not always reach full convergence. Thus, our results should be read as evidence of feasibility (existence of one-pass decoding) rather than a precise capacity bound.
Acknowledgments
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4F0002 and the agreement with Skoltech №139-10-2025-033.